OpenStudy (jtug6):

Find a taylor series and radius of convergence for f(x) centered at a = 1 where f(x) = (x^4)-(3x^2)+1

OpenStudy (jtug6):

So f(x) = (x^4)-(3x^2)+1 f'(x) = (4x^3)-(6x) f''(x) = (12x^2)-6 f'''(x) = 24x f''''(x) = 24 and f(1) = -1 f'(1) = -2 f''(1) = 12-6=6 f'''(1) = 24 f''''(1) = 24

OpenStudy (jtug6):

so taylor series is defined as \[\sum_{0}^{\infty} [f ^{n}(a)(x-a)^n]/n!\]

OpenStudy (jtug6):

so plugging that in i got -1 - 2(x-1)/1! + 3(x-1)^2 + 4(x-1)^3 + (x-1)^4 as the series

OpenStudy (jtug6):

@zepdrix some assistance? :D

OpenStudy (jtug6):

@ganeshie8 por favor? :-)

OpenStudy (jtug6):

@Kainui

OpenStudy (jtug6):

@ILovePuppiesLol

ILovePuppiesLol (ilovepuppieslol):

hi

ILovePuppiesLol (ilovepuppieslol):

whats that big symbol that looks like an E

OpenStudy (jtug6):

thatd be a sigma.

ILovePuppiesLol (ilovepuppieslol):

why are n exponents

OpenStudy (jtug6):

first ones actually the nth derivative, other ones an exponent

ILovePuppiesLol (ilovepuppieslol):

okey i officially am going to bed! i do not know this

OpenStudy (jtug6):

k

OpenStudy (loser66):

Taylor series is \(f(x-a) =\\ \color{red}{ f(a)} +\color{blue}{f'(a)} (x-a)+\color{green}{f''(a)} (x-a)^2+\color{orange}{f^{(3)}}(a) (x-a)^3+\color{purple}{f^{(4)}}(x-a)^4+f^{(5)}(x-a)^5+\cdots\) Your \(f^{(5)}(x)=0\) hence, the series stops at \(f^{(4)}\) Just plug the value in, you have \(f(x-1) = \color{red}{-1}+\color{blue}{(-2)}(x-1) +\color{green}{(6)}(x-1)^2 +\color{orange}{(24)}(x-1)^3+\color{purple}{(24)}(x-1)^4\) See what is going on??

OpenStudy (loser66):

For the radius, just plug x=1 in, all terms are gone but f(a) , hence radius is |-1|=1 . that means at a =1, inside the circle center 1, radius 1, the series converges.

OpenStudy (anonymous):

use polynomial division or manually manipulate it into form as a polynomial in \(x-1\) let \(x=u+1\) so $$x^4 -3x^2+1=(u+1)^4-3(u+1)^2+1=u^4+4u^3+3u^2-2u-1 $$ so $$f(x)=-1-2(x-1)+3(x-1)^2+4(x-1)^3+(x-1)^4 $$

OpenStudy (jtug6):

ah. so that essentially is the taylor series form? I dont need to represent it as some sort of sum because the derivative stopped after the fifth one? Meaning its finite?

OpenStudy (anonymous):

the radius of convergence for a polynomial is clearly infinite as the polynomial expression is a finite series. @Loser66 you forgot to divide by \(n!\) in your expression for the Taylor expansion

OpenStudy (anonymous):

yes, exactly @jtug6--the Taylor series gives a limit of successive polynomial approximations to a function using local data (derivatives of all orders); a polynomial itself though is already its best local polynomial approximation

OpenStudy (jtug6):

thank you!

OpenStudy (loser66):

:) yes, I did @oldrin.bataku